OpenStudy (anonymous):

HELP WILL FAN AND MEDAL!!! Solve each equation by factoring. 81x^2 = 49

OpenStudy (anonymous):

First step should be to move the 49 to the other side of the equation. This gives you 81x^2-49=0

OpenStudy (anonymous):

From there you should recognize that \[7^2=49\] and\[(9x)^2=81x^2\]

OpenStudy (anonymous):

Then you should use the formula \[a^2-b^2=(a-b)*(a+b)\]

OpenStudy (anonymous):

So that gives you \[(9x-7)(9x+7)\]

OpenStudy (anonymous):

Is that the answer?

OpenStudy (anonymous):

Yes that is the factor of that equation could you understand the steps

OpenStudy (anonymous):

A little... I suck at math.